A straight line passes through the points {x1, y1} = {4, 9} and {x2, y2} = {7, 5}. Find the equation of the line and consider the following statements. Which of these statements is true? (NB. More than one statement might be true. Select all statements that you think are true.)
1
Expert's answer
2021-08-23T13:11:07-0400
"\frac{x-4}{7-4}=\frac{y-9}{5-9}"

"4x+3y-43=0"